# Algorithmic and High-Frequency Trading Problems for Semi-Markov and Hawkes Jump-Diffusion Models

### v1

@@ 6 Conclusions and Future Recommendations
In this paper we introduced a new set of more general price processes that could be implemented in a number of algorithmic and HFT problems. The type of trading problems we focused on apply stochastic optimal control theory, where we chose an example acquisition problem from Cartea et al., (2015), as well as designing a similar liquidation problem, for illustrative purposes. Our results show that increasing the effect of the jump part (via a diffusion approximation) can significantly alter the optimal trading solution, thus inducing an agent to act differently in the presence of jumps in the asset price. The jump part in our price processes can be either a function of a Semi-Markov or Hawkes process, which have been proven to more accurately mimic LOB dynamics. We provide visuals for how the optimal solution changes based on the effects of jumps, as well as strategy simulations. The strategy simulations portray how the agent would act in different scenarios, based on different random paths of our price processes. It is clear from our analysis that the evolution of the acquisition and liquidation trading strategies, via simulations of the inventory and trading speed processes, can differ significantly based on this additional jump part. We would expect a similar effect across a wide range of trading problems using this type of SOC format, with some other examples including trading problems related to market-making, pairs trading or statistical arbitrage. In-depth overviews of these types of trading problems can be found in Cartea et al., (2015).
@@
Regarding future research, we would recommend applying similar logic to other types of trading problem setups not limited to the SOC framework used in this paper. For example, reinforcement or deep reinforcement learning (RL) has become very popular lately, and is becoming a state-of-the-art method for solving these types of trading problems. Gašperov and Kostanjčar, (2022) is one of the first to apply this type of innovative model in a deep RL setting under a Hawkes process pricing model, however, their model is just based on a jump case and no diffusion. Thus, we think this could be an interesting area to combine the two i.e., solve a deep RL trading problem under a jump-diffusion price process, where the jumps are either a function of a Semi-Markov or Hawkes process. We also believe it would be interesting to extend these price models to the Levy case, for cases where the LOB might follow a Markov process (since Semi-Markov and Hawkes processes are not Markov processes). Utilizing Levy processes would enable the use of a more general setting for price processes and would depend on the jump measure. However, we would like to point out that within the SOC framework for the acquisition and liquidation problems we studied in this paper, without our diffusion approximations, the problem would not be solvable under the numerical method we followed similarly to Cartea et al., (2015), as you would then be unable to follow the same dimension reduction method. An RL setup may be more adaptable to trading problems with these types of advanced price processes, as these kinds of limitations can often be alleviated then.
@@ Acknowledgments
We would like to thank MITACS and NSERC for research funding.

### v2

@@ 6 Conclusions and Future Recommendations
In this paper we introduced a new set of more general price processes that could be implemented in a number of algorithmic and HFT problems. The type of trading problems we focused on apply stochastic optimal control theory, where we chose an example acquisition problem from Cartea et al., (2015), as well as designing a similar liquidation problem, for illustrative purposes. These are are common execution trading problems for large institutional investors, who also often measure performance based on how well a large trade was executed. Our results show that increasing the effect of the jump part (via a diffusion approximation) can significantly alter the optimal trading solution, thus inducing an agent to act differently in the presence of jumps in the asset price. The jump part in our price processes can be either a function of a Semi-Markov or Hawkes process, which have been proven to more accurately mimic LOB dynamics. We provide visuals for how the optimal solution changes based on the effects of jumps, as well as strategy simulations. The strategy simulations portray how the agent would act in different scenarios, based on different random paths of our price processes. Average traded prices, as shown in Figures 5 and 11 for the acquisition and liquidation problems, respectively, are a heavily examined measure by institutions for the performance of trading algorithms focused on acquiring or liquidating large amounts of an asset. Thus, under our new models, the main implication is how a trading agent should optimally run specific execution trading algorithms like for when dealing with large positions, based on the more general pricing dynamics given in our Jump-Diffusion model. It is also clear from our analysis that the evolution of the acquisition and liquidation trading strategies, via simulations of the inventory and trading speed processes, can differ significantly based on this additional jump part, which are likely to appear in more volatile market regimes. We would expect a similar effect across a wide range of trading problems using this type of SOC format, with some other examples including trading problems related to market-making, pairs trading, or statistical arbitrage. In-depth descriptions of these types of trading problems in the SOC framework can be found in Cartea et al., (2015).
@@
Regarding future research, we suggest that the simulations in Section 5 be extended to encompass a wider range of market regimes. This could involve changing the parameter values in line with different market-regime rationals and then analyzing the performance of the acquisition and liquidation algorithms. We would also recommend applying similar logic to other types of trading problem setups not limited to the SOC framework used in this paper. For example, reinforcement or deep reinforcement learning (RL) has become very popular lately and is becoming a state-of-the-art method for solving these types of trading problems. Gašperov and Kostanjčar, (2022) is one of the first to apply this type of innovative model in a deep RL setting under a Hawkes process pricing model, however, their model is just based on a jump case and no diffusion. Thus, we think this could be an interesting area to combine the two i.e., solve a deep RL trading problem under a jump-diffusion price process, where the jumps are either a function of a Semi-Markov or Hawkes process. We also believe it would be interesting to extend these price models to the Levy case, for cases where the LOB might follow a Markov process (since Semi-Markov and Hawkes processes are not Markov processes). Utilizing Levy processes would enable the use of a more general setting for price processes and would depend on the jump measure. However, we would like to point out that within the SOC framework for the acquisition and liquidation problems we studied in this paper, without our diffusion approximations, the problem would not be solvable under the numerical method we followed similarly to Cartea et al., (2015), as you would then be unable to follow the same dimension reduction method. An RL framework may be better suited for trading problems involving advanced price processes, as it can help mitigate these limitations more effectively.
@@ Acknowledgments
The authors thank MITACS and NSERC for research funding, and the reviewers for their helpful and insightful comments.